OpenStudy (anonymous):

The elevation at the base of a ski hill is 350 m above sea level. A ski lift raises a skier (total mass = 72 kg, including equipment) to the top of the hill. If the skier’s gravitational potential energy relative to the base of the hill is now 9.2 × 105 J, what is the elevation at the top of the hill?

OpenStudy (jamesj):

What is the expression for gravitational potential energy, PE, in terms of mass \( m \), the gravitational acceleration \( g \) and the height \( h \) ?

OpenStudy (anonymous):

\[Eg = mgh\]

OpenStudy (jamesj):

Yes, \( PE = mgh \). Now suppose the skier is raised up an additional \( h \) from the base of the mountain, then you know that \[ PE = 9.2 \times 10^5 \] and \[ PE = mgh. \] You also know \( m \) and \( g \). Now solve for \( h \). Then the elevation of the mountain is \( 350 + h \).

OpenStudy (jamesj):

Make sense?

OpenStudy (anonymous):

yup, so that means 9.2 x 10^5 / 72 kg / 9.8 N/kg = h?

OpenStudy (jamesj):

looks right

OpenStudy (anonymous):

then I add 350 to the h to get the elevation?

OpenStudy (jamesj):

Yes, it must be, because h measures distance from the base, and the energy level 9.2x10^5 was specified as being " gravitational potential energy relative to the base of the hill "

OpenStudy (anonymous):

Oh i see now, so adding on the 350 makes it the height above sea level! Thank a lot dude!